Heights of men on a baseball team have a​ bell-shaped distribution with a mean of 183 cm183 cm and a standard deviation of 6 cm6 cm. Using the empirical​ rule, what is the approximate percentage of the men between the following​ values? a. 165165 cm and 201201 cm b. 177177 cm and 189189 cm
Answer #1

as we know k score =(X-mean)/std deviaation

a) therefore P(165<X<201)=P((165-183)/6<k<(201-183)/6)=P(-3<k<3)

as we know that 99.7% of values lies 3 std deviation from mean

hence  percentage of the men between 165 and 201 cm =99.7%

b) P(177<X<189)=P((177-183)/6<k<(189-183)/6)=P(-1<k<1)

as we know that 68.0% of values lies 1 std deviation from mean

hence  percentage of the men between 177 and 189 cm =68.0%